n ideal gas at 5 atm and 30°C. The ambient aa Pa. A valve is opened, half of the mass of the kconasure is en the valve is closed. If the final temperature in 5) A rigkd A rigid tank contains 5 kg of then the v 9as is allowed to escape, and the tank is -75 C, determine the final pressure inside the tank in atm
Answer #1

let n1 is the initial no of moles of the gas.

T1 = 30 C = 30 + 273 = 303 K


P1 = 5 atm

V1 = V

final no of moles of the gas, n2 = n1/2

T2 = -75 C = -75 + 273 = 198 K

V2 = V

P2 = ?


use,

P1*V1 = n1*R*T1 --(1)

P2*V2 = n2*R*T2 ---(2)

take equation(2)/equation(1)

P2*V2/(P1*V1) = n2*T2/(n1*T1)

P2 = P1*((n1/2)*198/(n1*303)

P2 = 5*198/(2*303)

= 1.634 atm